OpenStudy (lauluisa):

Write an equation in point-slope form of the line that passes through the given points (-3, 9) (10, -17)

OpenStudy (nurali):

first calculate slope (y2 - y1) / (x2 - x1) (x1,y1)= (-3, 9) (x2,y2)= (10, -17)

OpenStudy (nurali):

m=-2

OpenStudy (nurali):

secondly put value of slope i.e. m in the following eq. y = mx +b y = -2x + b

OpenStudy (nurali):

third substitute either point into the equation. either (-3, 9) or (10, -17) 9 = -2(-3) + b

OpenStudy (nurali):

Forth solve for b, which is the y-intercept of the line 9 = 6 + b b = 3

OpenStudy (nurali):

Fifth substitute b = 3, into the equation from step 2 y = -2x + 3 --- A A is your required equation :) hope you understand it !